OpenStudy (anonymous):

Write an equation of the line that passes through the point (3,0) and has the slope, m = -2.

OpenStudy (anonymous):

\[y-y1=m(x-x1)\]

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

You can use either..

OpenStudy (anonymous):

I suggest you to use @KlOwNlOvE equation.. :)

OpenStudy (anonymous):

x1=3 y1=0 m=-2